OpenStudy (anonymous):

How do you solve 5(b-4)=10b+5

hero (hero):

1. Distribute the 5: 5b - 20 = 10b + 5 Put the b variables on the same side, put the integers on the other side: 10 b - 5b = -(20 + 5) 5b = -25 Solve for b: b = -25/5 b = -5

OpenStudy (anonymous):

Thanks